OpenStudy (anonymous):

Perform the indicated operation. 4/7x^2 - 3/2x^2

OpenStudy (ivancsc1996):

You can simplify to this:\[\frac{ 4 }{ 7x^{2} }-\frac{ 3 }{ 2x^{2} }=\frac{ 1 }{ x^{2} }(\frac{ 4 }{ 7 }-\frac{ 3 }{ 2 })=\frac{ 1 }{ x^{2} }(-\frac{ 13 }{ 14 })=-\frac{ 13 }{ 14x^{2} }\]